in a random sample of six microwave ovens, the mean repair cost was $60.00 and the standard deviation was $11.00

Round to two decimal places as needed) (Round to two decimal places as needed.) C. It can be said that 99% of microwaves have a repair cost between the bounds of the c interval DOLL
Answer #1

Solution:- Given that mean = 60.00, standard deviation = 11, n = 6

df = n-1 = 6-1 = 5

t-value = 4.032

99% confidence interval for the population mean = X +/- t*s/sqrt(n)
= 60 +/- 4.032*11/sqrt(6)
= 41.89 , 78.11

The margin of error is

=> t*s/sqrt(n) = 4.032*11/sqrt(6) = 18.11

=> option D.

table :

t Table cum. prob one-tai 0.50 0.25 0.20 0.15 0.10 0.05 0.025 0.01 0.005 two-tails 1.00 0.50 0.40 0.30 0.20 0.10 0.05 0.02 0.01 6 1 0.000 1.000 .376 1.963 3.078 6.3147 31.82 63.66 2 0.000 0.816 .0611.386 1.886 2.920 4.303 6.965 9.925 3 0.000 0.765 0.978 1.250 638 2.353 3.182 4.541 5.84 4 0.000 0.74 0.941.190 1.533 2.132 2.776 3.747 4.604 5 0.000 0.727 0.920 1.156 1.476 2.015 2.573.365 4.032